OpenStudy (ttop0816):

tan110-tan80/1+tan110*tan50=??tan2.2° tan60° tan160° tan550°

OpenStudy (ttop0816):

tan2.2°

OpenStudy (***[isuru]***):

can u find the same pattern in ur question too ?

OpenStudy (***[isuru]***):

\[\tan ( a- b ) = \frac{ \tan \ a - \tan b }{ 1 + \tan \ a \ \tan \ b }\]

OpenStudy (ttop0816):

yes i have the same pattern!

OpenStudy (***[isuru]***):

\[\tan ( 110 - 50 ) = \frac{ \tan 110 - \tan 50 }{ 1 + \tan 110 \tan50}\]

OpenStudy (***[isuru]***):

so... u can find ur answer now :)

OpenStudy (ttop0816):

then would it be tan60? (:

OpenStudy (***[isuru]***):

yep:)

OpenStudy (ttop0816):

Thank you !!

OpenStudy (***[isuru]***):

u r welcome!

OpenStudy (gudden):

@***[ISURU]*** .. u are answering..! :p